OpenStudy (anonymous):

A particle in a circle at 200rpm. If the frequency is increased to 400rpm in 2 min at a uniform rate the magnitude of angular acceleration is?

OpenStudy (anonymous):

Final angular velocity = Initial angular velocity + angular accl .* time Let x be the time, 400 = 200 + 2*x , x = 100 revolutions/min^2

OpenStudy (anonymous):

@kartiksriramk wat we have to find in ur eq

OpenStudy (anonymous):

We have to find the angular acceleration.

OpenStudy (anonymous):

@kartiksriramk wat should i do now

OpenStudy (anonymous):

The angular acceleration is the answer.That all.

OpenStudy (anonymous):

i have to solve for x

OpenStudy (anonymous):

400 = 200 + 2*x

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

x =100

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

The option are pie/100 pie/200 pie/36 pie/18

OpenStudy (experimentx):

just find the angular acceleration. final - initial / time ... and change those all into radians.

OpenStudy (anonymous):

well then you have to convert them into radians/second^2.

OpenStudy (anonymous):

It is pi/18. Multiply the answer by 2pi/60^2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i gor it

OpenStudy (anonymous):

pie/18

OpenStudy (anonymous):

thxxx